how do I factor and then solve using the principal of zero?
2x^2 +20 = x^2 + 9x
do I
add -2x^2 to both sides
20 = -x^2 + 9x
add -20 both sides
0 = - x^2 + 9x -20
is this right?
then what?

Found 3 solutions by Alan3354, Grinnell, mananth:
Answer by Alan3354(69443)

how do I factor and then solve using the principal of zero?
2x^2 +20 = x^2 + 9x
do I
add -2x^2 to both sides
20 = -x^2 + 9x
add -20 both sides
0 = - x^2 + 9x -20
is this right?
then what?
-------------
That's right. Make it easier by multiplying by -1
x^2 - 9x + 20 = 0
(x-5)*(x-4) = 0
x = 5
x = 4

Answer by Grinnell(63)

You made an error, we all do, don't worry...
OK 2x^2+20=x^2+9x
add negative x^2 to both sides
x^2+20=9x
X^2-9x+20=0
(x-4)(x-5)=0
solutions: x=4, and x=5 MINUS 9 X WAS ADDED TO BOTH SIDES, THIS YOU DID NOT DO
GET SOME REST!

Answer by mananth(16949)

how do I factor and then solve using the principal of zero?
2x^2 +20 = x^2 + 9x
do I
add -2x^2 to both sides
20 = -x^2 + 9x
add -20 both sides
0 = - x^2 + 9x -20
is this right?
then what?
-(0) = -(- x^2 + 9x -20)
Make the expressions -ve on both sides
x^-9x+20=0
x^2-5x-4x+20
x(x-5)-4(x-5)=0
(x-5)(x-4)=0
x=5 or4